OpenStudy (anonymous):

i can youtube this one unless anyone knows this one off the top of their head and could show work If 2x + 3y = 10, then the y-intercept of the graph of this equation is

OpenStudy (anonymous):

2x + 3y = 10 ===> y= -2/3 +10/3 ===> the y-intercept of the graph of this equation is 10/3

OpenStudy (chaise):

2x + 3y = 10 Rearrange this to put it into point-slope formula. 2x + 3y = 10 2x-10 = -3y (2/-3)x-10/-3=y y = (2x/-3)+(10/3) Want me to use equation editor?

OpenStudy (anonymous):

2x + 3y = 10 ===> y= -2/3x +10/3 ===> the y-intercept of the graph of this equation is 10/3

OpenStudy (chaise):

Raheen is right - y=mx+b, m is your slope, b is your y intercept. Your y intercept is 10/3

OpenStudy (anonymous):

just checked it. i was looking for y=mx+b now i know how to do these again.. Thanks Chaise and Raheen. I understand how it is 10/3

OpenStudy (chaise):

Point-slope formula is y=mx+b. Should have explained, sorry.

OpenStudy (anonymous):

you are welcome.

OpenStudy (anonymous):

yes as chaise pointed out.